One leg of a right triangle is 2 inches smaller than the hypotenuse. The other leg is 5 inches long. What are the lengths of the unknown leg and hypotenuse?
Not sure if I set it up right.
(x+2)^2 + 5^2 = x^2

Answer by checkley71(8403)

LET X BE THE HYPOTENUSE.
ONE SIDE=(X-2)
THE OTHER SIDE=5
(X-2)^2+5^2=X^2
X^2-4X+4+25=X^2
-4X+29=0
-4X=-29
X=-29/-4
X=7.25 ANSWER FOR THE LENGTH OF THE HYPOTENUSE.
7.25-2=5.25 ANSWER FOR THE LENGTH OF THE UNKNOWN LEG.
PROOF
5.25^2+25=7.25^2
27.56+25=52.56
52.56=52.56